OpenStudy (anonymous):

(03.02) If g(x) = 3(x - 2), find the value of x if g(x) = 6.

OpenStudy (anonymous):

I think x = 4

OpenStudy (michele_laino):

@AhhhhhplzzzzzHelp that's right!

OpenStudy (anonymous):

i think its 12 because the 6 would substitute the x in the equation right?

OpenStudy (michele_laino):

@Catfish15 please you have to solve the equation below: 3(x-2)=6, please, divide by 3, that equation, and write your result

OpenStudy (anonymous):

i'm not sure if that is the correct way to do it

OpenStudy (michele_laino):

why?

OpenStudy (anonymous):

@StudyGurl14 help please

OpenStudy (anonymous):

our already given the value 6 their just asking to input it into the equation, but i could be wrong

OpenStudy (michele_laino):

@Catfish15 please, substitute x=4 into the left side of your equation, and tell me what you get

OpenStudy (anonymous):

that doesn't make any sense why would you do that

OpenStudy (michele_laino):

@Catfish15 because that is the only way to verify an equation, after I found the solutions

OpenStudy (anonymous):

because when you put 4 in for x and 6 in for g(x) it make 6=6

OpenStudy (anonymous):

f(6)=3(4-2) f(6)=12-6 f(6)=6

OpenStudy (anonymous):

i got it now thank@AhhhhhplzzzzzHelp

OpenStudy (anonymous):

yeeep

OpenStudy (anonymous):

how do i send metals

OpenStudy (anonymous):

by press best response